OpenStudy (anonymous):

Complete the square: Y:x^2-8x+21 Someone plz help.

OpenStudy (radar):

1. Insure the coefficient of the squared term is 1, this case it is. 2. Compute 1/2 the value of the coefficient of the x term. 1/2 of 8 is 4. 3. Square this. Get 16 Now add this to the equation: \[x ^{2}-8x + 16 +21= 16\] 4. Subtract 21 from both sides getting \[x ^{2}-4x + 16 = -5\] 5. Express the left as a square \[(x - 4)^{2}=-5\] 6. Take the square root of both sides. \[x - 4 = \sqrt{-5}\] 7. Solve getting\[x = 4\pm \sqrt{-5}\] There you are with two comples roots.

OpenStudy (radar):

Please correct step 4. to read as:\[x ^{2}-8x +16 = -5\]